OpenStudy (anonymous):

304/100 in a fraction in lowest term

OpenStudy (anonymous):

76/25= 3.04

OpenStudy (anonymous):

97/25=3.04

OpenStudy (anonymous):

97/25=3.04

OpenStudy (anonymous):

hmm i wonder if it is 97/25?

OpenStudy (anonymous):

or 96/25

OpenStudy (anonymous):

lol it is if she asked 384/100